Circular No. 8276

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

S/2003 J 22 Further to IAUC 8138, the discovery of a new satellite of Jupiter has

been reported on CCD images taken with the 3.6-m Canada-France-Hawaii Telescope and the 2.2-m University of Hawaii reflector by teams led by S. S. Sheppard and by B. Gladman. Astrometry, preliminary orbital elements by B. G. Marsden, and an ephemeris appear on MPEC 2004-B41 (e = 0.23, i = 151^{o}, P = 640 days, H = 16.4).

SUPERNOVAE 2004M, 2004N, AND 2004O A. Udalski, Warsaw University Observatory, on behalf of the OGLE-

III survey team, reports the discovery with the OGLE/NOOS transient detection system (Udalski 2003, Acta Astron. 53, 291) of three apparent supernovae near galaxies located behind the Magellanic-Cloud halo.

SN  2004 UT  α_{2000}  δ_{2000}  I  Offset
2004M Jan. 6.1 0^{h}35^{m}59.^{s}26 −74^{o}48′02′′.0 19.5 10′′.7 W, 3′′.1 S
2004N Jan. 21.2 5 35 20.37 −69 27 18.3 18.0 4′′.4 N
2004O Jan. 24.2 4 43 44.82 −70 28 33.8 20.0 3′′.6 W, 1′′.6 N

Additional I magnitudes from OGLE-III images: SN 2004M, 2003 Dec. 19.0 UT, [22:; 22.1, 19.8; 25.0, 19.5; 28.0, 19.7; 31.0, 19.7; 2004 Jan. 3.0, 19.8; 12.0, 20.2; 15.0, 20.2; 17.1, 20.0; 19.0, 20.2; 21.0, 20.2; 23.1, 20.0. SN 2004N, 2004 Jan. 14.3, [22:; 16.3, 19.8; 17.3, 19.4; 19.2, 18.8; 23.2, 18.0; 25.2, 17.9. SN 2004O, 2004 Jan. 14.2, [22:; 16.2, 21.5; 18.3, 20.4; 20.2, 20.0; 22.2, 20.2; 24.2, 20.1. Finding charts can be found at

http://ogle.astrouw.edu.pl/ogle3/ews/NOOS/noos.html.

COMETS C/2003 B2, C/2003 C1, AND C/2003 D1 (SOHO) Further to IAUC 8269, K. Battams reports measures of additional

Kreutz sungrazing comets found on SOHO C3 website images by M. Oates (C/2003 B2) and X. Leprette (C/2003 C1 and C/2003 D1). Reduced positions and orbital elements by B. G. Marsden appear on MPEC 2004-B34.

Comet  2003 UT  α_{2000}  δ_{2000}
C/2003 B2 Jan. 10.821  20^{h}15.^{m}4  −22^{o}05′
C/2003 C1 Feb. 3.854  21 24.7  −18 14
C/2003 D1  28.321  23 03.2  − 9 00
2004 January 25  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green